Diazepam is used as a skeletal muscle relaxant and as a sedative administered before surgery, as well as a treatment for anxiety. For children above the age of 6 months who suffer from anxiety disorders, diazepam is administered at a reduced dosage. Suppose a child with a body surface area (BSA) of 0.8 m2 is to receive a dose of diazepam divided into three administrations. Using the BSA rule and assuming that an adult dose of diazepam is 12 mg, what are the daily dose and the dose per administration?
Answer #1

Child dose = BSA/ 1.73 * adult dose

= 0.8 / 1.73 * 12

= 0.46 * 12

= 5.52

Therefore the daily dosage for the child is 5.52mg.

One dose per administration = 5.52 / 3

= 1.84 mg per dose for the child .